OpenStudy (anonymous):

Graph y = |x + 2| and give domain and range

OpenStudy (solomonzelman):

Rules of \(\large\color{black}{ \rm shifts }\) from \(\large\color{black}{ \rm f(x) }\) to \(\large\color{black}{ \rm g(x) }\). \(\large\color{black}{ \rm f(x)=\left| x \right| ~~~~~~~~~\rm{\Longrightarrow}~~~~~~~~\rm g(x)=\left| x \color{blue}{ -~\rm{c} }\right| }\) \(\large\color{blue}{ ~\rm {c} }\) units to the \(\normalsize\color{blue}{ \rm right }\). \(\large\color{black}{ \rm f(x)=\left| x \right| ~~~~~~~~~\rm{\Longrightarrow}~~~~~~~~\rm g(x)=\left| x \color{blue}{ +~\rm{c} }\right| }\) \(\large\color{blue}{ ~\rm {c} }\) units to the \(\normalsize\color{blue}{ \rm left }\). \(\large\color{black}{ \rm f(x)=\left| x \right| ~~~~~~~~~\rm{\Longrightarrow}~~~~~~~~\rm g(x)=\left| x \right| \color{blue}{ +~\rm{c} }}\) \(\large\color{blue}{ ~\rm {c} }\) units \(\normalsize\color{blue}{ \rm up }\). \(\large\color{black}{ \rm f(x)=\left| x \right| ~~~~~~~~~\rm{\Longrightarrow}~~~~~~~~\rm g(x)=\left| x \right| \color{blue}{ -~\rm{c} }}\) \(\large\color{blue}{ ~\rm{c} }\) units \(\normalsize\color{blue}{ \rm down }\). So you know how your function is shifted, right?

OpenStudy (solomonzelman):

For the domain, we don't see any restrictions, do we? And for the range, we know an absolute value has to be equal to 0 or to a greater number, but not to a negative.

OpenStudy (anonymous):

No, not really.

OpenStudy (solomonzelman):

you see you added 2 into the absolute value, so look at the rule number 2. What is done to your graph?

OpenStudy (anonymous):

Im completely confused

OpenStudy (solomonzelman):

y = |x `+ 2`| see the gray part?

OpenStudy (solomonzelman):

Look at rule, 2 where you add +C into the absolute value.